Charlotte was asked to determine the concentration of ammonia in a commercially available cloudy ammonia solution used for cleaning. First she pipetted 25.00mL of the cloudy ammonia solution into a 250.00mL conical flask. 50.00mL of 0.100M HCl(aq) was added to the conical flask which immediately reacted with the ammonia in solution. The excess, (unreacted) HCl was then titrated with 0.050M Na2CO3(aq). 21.50mL of Na2CO3(aq) was required. Help Charlotte out by calculating the concentration of the ammonia in the cloudy ammonia solution.
1
Expert's answer
2020-07-09T14:56:39-0400

2HCl(aq)+Na2CO3(aq)→2NaCl(aq)+CO2(g)+H2O(l)

acid + carbonate→ salt + carbondioxide + water

moles = concentration (mol L-1) × volume (L)

n(Na2CO3(aq)) = c × V

c(Na2CO3(aq)) = 0.050 mol L-1

V(Na2CO3(aq)) = 21.50 mL = 21.50 × 10-3 L

n(Na2CO3(aq)) = 0.050 × 21.50 × 10-3 = 1.075 × 10-3 mol

From the balanced chemical equation, 1 mole Na2CO3 react with 2 moles of HCl

So, 1.075 × 10-3 mole Na2CO3 reacted with 2 × 1.075 × 10-3 moles HCl

n(HCltitrated) = 2 × 1.075 x 10-3 = 2.150 × 10-3 mol

The amount of HCl that was added to the cloudy ammonia solution in excess was :

n(HClexccess) = 2.150 × 10-3 mol

n(HCltotal added) = c × V

c(HCltotal added) = 0.100 mol L-1

V(HCltotal added) = 50.00 mL = 50.00 × 10-3 L

n(HCltotal added) = 0.100 × 50.00 × 10-3 = 5.00 × 10-3 mol

n(HCltitrated) + n(HClreacted with ammonia) = n(HCltotal added)

n(HCltotal added) = 5.00 × 10-3 mol

n(HCltitrated) = 2.150 × 10-3 mol

2.150 × 10-3 + n(HClreacted with ammonia) = 5.00 × 10-3

n(HClreacted with ammonia) = 5.00 × 10-3 - 2.150 × 10-3 = 2.85 × 10-3 mol

NH3(aq) + HCl(aq) → NH4Cl(aq)

From the equation, 1 mol HCl reacts with 1 mol NH3

So, 2.85 × 10-3 mol HCl had reacted with 2.85 × 10-3 mol NH3 in the cloudy ammonia solution

concentration (mol L-1) = moles ÷ volume (L)

c(NH3(aq)) = n(NH3(aq)) ÷ V(NH3(aq))

n(NH3(aq)) = 2.85 × 10-3 mol (moles of NH3 that reacted with HCl)

V(NH3(aq)) = 25.00 mL = 25.00 × 10-3 L (volume of ammonia solution that reacted with HCl)

c(NH3(aq)) = 2.85 × 10-3 ÷ 25.00 × 10-3 = 0.114 mol L-1